Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Sposito - 16:01 Friday 01 July 2022 (56326) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned TCS activity NI CH centering measurements performed by C. Taranto. We found an issue with TCSPicoNI,  we tried to stop and restart the process and to manually switch off and on the driver but we still have problems. 13:50 UTC I went into the TCS room in order to swap the NI and WI picomotors driver in order to see if it was the problem an investigation is ongoing, afternoon operator was informed about all situation.
Activity still in progress.

No DMS Events to report

Sub-system reports

10:47 UTC TCSPicoNI not working properly, so I stopped and restarted the process.
around 12:05 UTC F. Carbognani went into the TCS room to manually reset the PC unit that controls the Picomotors of the TCS benches in order to fix a problem.

Comments to this report:
carbognani, gherardini, sposito, taranta - 18:59 Friday 01 July 2022 (56328) Print this report

As far as the problems reported to move the NI_HM_tx/ty picomotors we tryied:

  • stop/restart the TCSPicoNI server: no effects
  • reset (OFF/ON) the controller: no effects
  • swap WI and NI controllers: no effects

From the software/config point of view:

  • the ActuatorApp corresponding software (version v5r1p1) is unchanged from 2018
  • TCSPicoNI server config file is unchanged from 2016

We tryied to send command from a terminal via cm send just to exclude any problem from the Python client side but we had the same result.

The behaviour is that while moving both tx and ty channels in Forward direction, logs from the server report the move is performed but from the HWS we see no move at all,  while moving in Backward direction the move is reported as not performed by the controller. Correspondig logs below.

2022-07-01-12h23m11-UTC>INFO...-MoveRel NI_HM_ty F 5500.000000 command received from position -53337.000000
2022-07-01-12h23m15-UTC>INFO...-Memory Used increase, cur. 3884.00(KB), inc. 24.00(KB)
2022-07-01-12h23m19-UTC>INFO...-Ping command received
2022-07-01-12h23m40-UTC>INFO...-Ping command received
2022-07-01-12h23m51-UTC>INFO...-Ping command received
2022-07-01-12h24m09-UTC>INFO...-main> Stopped at position = -47837.000000

2022-07-01-10h30m06-UTC>INFO...-MoveRel NI_HM_tx B 5500.000000 command received from position -50365.000000
2022-07-01-10h30m08-UTC>INFO...-main> Stopped at position = -50365.000000

Check to hear the picomotor  noise from outside the TCS bench enclousure were inconclusive. Inspection on the bench itself  seems needed., 

Search Help